OpenStudy (anonymous):

solve using the addition principal 2x+8 is less than or equal to x+9

OpenStudy (anonymous):

Imagine like you had an \(=\) sign instead of a \(\le\) sign and solve for \(x\) just as you would normally EXCEPT for the fact that you must flip the sign if you divide both sides by a negative number (which you don't necessarily need to worry about for this one). Show me how far you've gotten so far.

OpenStudy (anonymous):

I havent gotten anywhere I an completely dumbfounded when it comes to algebra

OpenStudy (anonymous):

2x + 8 <= x+9 is the given piece of information. Now, you need to gather terms with x to one side and constants to the other. If you do that right, you get 2x - x <= 9 -8 Now, simple subtraction will get you: x <= 1